# DiffCut: Catalyzing Zero-Shot Semantic Segmentation with Diffusion Features and Recursive Normalized Cut

### v2

@@ Abstract
Foundation models have emerged as powerful tools across various domains including language, vision, and multimodal tasks. While prior works have addressed unsupervised image segmentation, they significantly lag behind supervised models. In this paper, we use a diffusion UNet encoder as a foundation vision encoder and introduce DiffCut, an unsupervised zero-shot segmentation method that solely harnesses the output features from the final self-attention block. Through extensive experimentation, we demonstrate that the utilization of these diffusion features in a graph based segmentation algorithm, significantly outperforms previous state-of-the-art methods on zero-shot segmentation.  Specifically, we leverage a recursive Normalized Cut algorithm that softly regulates the granularity of detected objects and produces well-defined segmentation maps that precisely capture intricate image details. Our work highlights the remarkably accurate semantic knowledge embedded within diffusion UNet encoders that could then serve as foundation vision encoders for downstream tasks. Project page: https://diffcut-segmentation.github.io
@@ 1 Introduction
In this work, we introduce DiffCut, a new method for zero-shot image segmentation which solely harnesses the encoder features of a pre-trained diffusion model in a recursive graph partitioning algorithm  to produce fine-grained segmentation maps. Importantly, our method does not require any label from downstream segmentation datasets and its backbone has not been pre-trained on dense pixel annotations such as SAM [27].  We observe in Fig. 1 that DiffCut produces sharp segments that nicely outline object boundaries.  In comparison with the recent state-of-the-art unsupervised zero-shot segmentation method DiffSeg [1], the segments yielded by DiffCut, are better aligned with the semantic visual concepts, e.g. DiffCut is able to uncover the urban area as well as the boats in the middle row image. Our main contributions are as follows:
  • We leverage the features from the final self-attention block of a diffusion UNet encoder, for the task of unsupervised image segmentation. In this context, we demonstrate that exploiting the inner patch-level alignment yields superior performance compared to merging self-attention maps as done in DiffSeg [1].
  • Compared to existing graph based object localization methods e.g. TokenCut or MaskCut [30, 32], we push further and take advantage of a recursive Normalized Cut algorithm to generate dense segmentation maps. Via a partitioning threshold, the method is able to regulate the granularity of detected objects and consequently adapt the number of segments to the visual content.
@@ 3.1 Features Extraction
An input image  is encoded into a latent via the VQ-encoder of the latent diffusion model and a small amount of gaussian noise is added to it (not shown in Fig. 2). The obtained latent is passed to the diffusion UNet encoder, from which we only extract the output features, denoted $\hat{z}$, from its last self-attention block.  This choice design has several motivations:
@@
In contrast to several methods that harness cross-attention modules for localization or segmentation tasks [39, 40, 41, 42], we deliberately choose not to depend on this mechanism. The accuracy of segmentation maps generated via attention modules heavily relies on the quality of the textual input which often requires an automatic captioning model combined with a meticulous prompt design to reach competitive performance. Besides being constrained by the maximum number of input tokens, such approach is proved to be inaccurate in the presence of cohyponyms [64] and is prone to neglect subject tokens as the number of objects to detect becomes large [65]. The localization and segmentation capacity with a single forward inference is then constrained by the performance of the captioning model and the attention modules themselves. Exploiting only the intermediate diffusion features alleviate the computational cost of an additional captioning model and do not necessitate multiple forward inferences.
@@
Previous works [66, 67, 37] have shown that diffusion features provide precise semantic information shared across objects from different domains. Building on this observation, we hypothesize that the pyramidal architecture of the UNet encoder capture semantically rich image representations that are well-suited for zero-shot vision tasks. To validate this assumption, we exhibit  the semantic coherence emerging in the UNet encoder, evidenced by a remarkable patch level alignment in the output features of the final self-attention block. We in fact demonstrate  that these features are sufficient to reach state-of-the-art zero-shot segmentation performance.
@@ 3.2 Recursive Feature Clustering
to find the eigenvector $x$ corresponding to the second smallest eigenvalue. In the ideal case, the clustering solution only takes two discrete values.  Since the solution of Eq. 2 is a continuous relaxation of the initial problem, $x$ contains continuous values and a splitting point has to be determined to partition it.  To find the optimal partition, we examine $l$ evenly spaced points in $x$ and select the one resulting in the minimum NCut value.
@@
Deriving from our observation on the patch-level alignment of our chosen diffusion features in Fig. 4, we assume that the normalized cut algorithm would provide sharp segments, each corresponding to a precise semantic concept as distinct objects would manifest as weakly connected components in a patch similarity matrix. Following this intuition, we construct an affinity matrix $\bm{W}$, by computing the cosine similarity between each pair of patches.  In addition, since NCut criterion assesses both the overall dissimilarity among different segments and the similarity within each segment, we opt to emphasize inter-segments dissimilarity by exponentiating each element with a positive integer value $\alpha$:
@@
Classical spectral clustering [68] requires setting a pre-defined number of clusters to partition the graph, which is a significant constraint in the context of zero-shot image segmentation where no prior knowledge on the visual content is available. We therefore adopt a recursive graph partitioning [55], which adapts the number of uncovered segments to the visual content via a threshold, denoted $\tau$, on the maximum partitioning cost.  This hyperparameter stops the recursive partitioning of a segment when its NCut value exceeds it and thereby regulates the granularity of detected objects. We demonstrate that our soft thresholding process detailed above enhances the robustness of the method which delivers competitive performance across a wide range of $\tau$ values. This recursive clustering process is summarized in Supplementary A.
@@ 4 Experiments
Following existing works in image segmentation [21, 53, 52, 18], we use the following datasets for evaluation: a) Pascal VOC [70] (20 foreground classes), b) Pascal Context [71] (59 foreground classes), c) COCO-Object [72] (80 foreground classes), d) COCO-Stuff-27 merges the 80 things and 91 stuff categories in COCO-stuff into 27 mid-level categories, e) Cityscapes [73] (27 foreground classes) and f) ADE20K (150 foreground classes) [74]. An extra background class is considered in Pascal VOC, Pascal Context, and COCO-Object. We ignore their training sets and directly evaluate our method on the original validation sets, at the exception of COCO for which we evaluate on the validation split curated by prior works [21, 19].
@@
For all datasets, we report the mean intersection over union (mIoU), the most popular evaluation metric for semantic segmentation. Because our method does not provide a semantic label, we use the Hungarian matching algorithm [75] to assign predicted masks to a ground truth mask. For datasets including a background class, we perform a many-to-one matching to the background label. As in [1], we emphasize unsupervised adaptation (UA), language dependency (LD), and auxiliary image (AX). UA means that the specific method requires unsupervised training on the target dataset. This is common in the unsupervised segmentation literature. Methods without the UA requirement are considered zero-shot. LD means that the method requires text input, such as a descriptive sentence for the image, to facilitate segmentation. AX means that the method requires an additional pool of reference images or synthetic images.
@@
DiffCut builds on SSD-1B [35], a distilled version of Stable Diffusion XL [34]. The model takes an empty string as input and we set the timestep for denoising to $t=50$. To ensure a fair comparison when evaluating our method against baselines, we set a unique value for $\tau$ and $\alpha$ across all datasets, equal to $0.5$ and $10$ respectively.  Following previous works, we make use of PAMR [69] to refine our segmentation masks. Our method runs on a single NVIDIA TITAN RTX (24GB) with input images of size $1024\times 1024$ and can segment an image in one second.
@@ 4.1 Results on Zero-shot Segmentation
Tab. 1 reports the mIoU score for each baseline across the 6 benchmarks. Note that the numbers shown for COCO-Stuff and Cityscapes are taken from [1].  We complete ReCo [53] and MaskCLIP [52] scores with the results obtained in [50].  Other numbers are taken from [18].  We also note that DiffSeg tunes the sensible merging hyperparameter on a subset of images from the training set from the respective datasets. For a fair comparison, we evaluate the method fixing it to 1, as recommended in the original paper, and refine the obtained masks with PAMR. This baseline is denoted DiffSeg.
@@
Model LD AX UA VOC Context COCO-Object COCO-Stuff-27 Cityscapes ADE20K Extra-Training    IIC [19] 9.8 - - 6.7 6.4 -    MDC [76] - - - 9.8 7.1 -    PiCIE [21] - - - 13.8 12.3 -    PiCIE+H [21] - - - 14.4 -    STEGO [20] - - - 28.2 21.0 -    ACSeg [18] 53.9 - - 28.1 - - Training-free    ReCO [53] 25.1 19.9 15.7 26.3 19.3 11.2    MaskCLIP [52] 38.8 23.6 20.6 19.6 10.0 9.8    MaskCut $(k=5)$ [32] 53.8 43.4 30.1 41.7 18.7 35.7    DiffSeg [1] - - - 43.6 21.2 -    DiffSeg 49.8 48.8 23.2 44.2 16.8 37.7    DiffCut (Ours) 65.2 56.5 34.1 49.1 30.6 44.3
With our set of default hyperparameters, DiffCut significantly outperforms all other baselines despite not relying on language dependency, auxiliary images or unsupervised adaptation. On average, our method achieves a gain of +7.3 mIoU over the second best baseline. Notably, DiffCut exceeds MaskCut with an average improvement of +9.4 mIoU. Additionally, it outperforms the previous state-of-the-art method in unsupervised segmentation, DiffSeg, by +5.5 mIoU on COCO-Stuff and +9.4 mIoU on Cityscapes. The superiority of DiffCut over these two methods demonstrates our two key contributions: the high quality of our visual features for semantic segmentation and the flexibility of the recursive NCut algorithm in adjusting the number of segments according to the visual content of each image. The effectiveness of our method is further corroborated by our qualitative results shown in Fig. 1. In comparison to DiffSeg, DiffCut provides finely detailed segmentation maps that more closely align with semantic concepts. Additional examples can be found in Supplementary L.
We note here that, as the granularity of annotations varies across target datasets, our fixed set of hyperparameters can not be in the optimal regime on each of them. Therefore, relaxing the condition on prior knowledge about the target dataset, we report in Supplementary G results of DiffCut where $\tau$ is loosely tuned using a small subset of annotated images from the target training split.
@@ 4.2 Semantic Coherence in Vision Encoders
As in [79], we collect patch representations from various vision encoders and store their corresponding target classes using the segmentation labels. Given $\hat{z}_{i}=\mathcal{E}(\mathbf{x_{1}})_{i}\in\mathbb{R}^{D_{v}}$ and $\hat{z}_{j}=\mathcal{E}(\mathbf{x_{2}})_{j}\in\mathbb{R}^{D_{v}}$, the patch representations of images $\mathbf{x_{1}}$ and $\mathbf{x_{2}}$ at respectively index $i$ and $j$, we compute their cosine similarity  and use this score as a binary classifier to predict if the two patches belong to the same class. Given $l(\mathbf{x_{1}})_{i}$ and $l(\mathbf{x_{2}})_{j}$, the labels associated to the patches, if $l(\mathbf{x_{1}})_{i,j}=l(\mathbf{x_{2}})_{p,q}$, the target value for binary classification is 1, else 0. We present in Fig. 3 the ROC curve and AUC score for our candidate models. We observe that SSD-1B UNet encoder [35] demonstrates a greater patch-level alignment than any other candidate model with an AUC score of 0.83, even surpassing DINOv2 [5]. We further exhibit the outstanding alignment between patch representations associated to semantically similar concepts with qualitative results in Fig. 4. We provide additional qualitative examples patch-level alignment in Supplementary K.
@@
A potential rationale for this observation lies in the superior semantic information retention of a diffusion model compared to alternative backbones, attributed to its inherent capacity to set a structural image layout, internally acquired during the training phase. These results provides insight into the strong clustering results presented in previous section, as improved semantic coherence suggests that patches belonging to the same object are more effectively clustered.
@@ 4.3 Ablation study
DiffSeg proposes their own clustering algorithm based on a self-attention map merging process. As the original implementation uses a different diffusion backbone as ours,  we validate the benefit of our method by swapping the original SDv1.4 with our stronger SSD-1B. For a fair comparison between methods, we use the default set of hyperparameters recommended in [1] and set the default merging threshold of DiffSeg to $0.5$ for all datasets. Tab. 2 clearly validates the superiority of using rich semantic features in a recursive graph partitioning algorithm over the self-attention merging mechanism of DiffSeg. Qualitative results shown in Fig. 1 further display the edge of DiffCut in uncovering semantic clusters. Shown results do not make use of the mask refinement module, explaining the gap with Tab. 1.
@@
In DiffCut, the hyperparameter $\tau$ corresponds to the maximum graph partitioning cost allowed.  In contrast, classical spectral clustering requires to explicitly set the number of segments to be found in the graph. To validate the benefit of the recursive approach over spectral clustering, we introduce a simple yet effective baseline dubbed AutoSC. [80] proposes a heuristic that estimates the number of connected components in a graph with the largest relative-eigen-gap between its Laplacian eigen-values. The larger the gap, the more confident the heuristic. In our context, the index of the eigen-value that maximizes this quantity can be interpreted as the number of clusters in an image. Thus, we define a set of exponents $\{1,5,10,15\}$ and determine the value $\alpha$ in this set such that its element-wise exponentiation of matrix $\bm{A}$ yields the largest Laplacian relative-eigen-gap. Then, we use the index of the eigen-value maximizing the gap as the number of clusters in a $k$-way spectral clustering performed with the algorithm proposed in [81]. As shown in Tab. 2, DiffCut consistently outperforms AutoSC on all datasets, with a gain up to $+3.5$ on ADE20K, at the exception of COCO-Stuff where the latter yields slightly better results. Noting that AutoSC is already a state-of-the-art baseline  on most benchmarks, this experiment confirms the relevance of the recursive Normalized Cut to uncover arbitrary numbers of segments.

### v3

@@ Abstract
Foundation models have emerged as powerful tools across various domains including language, vision, and multimodal tasks. While prior works have addressed unsupervised semantic segmentation, they significantly lag behind supervised models. In this paper, we use a diffusion UNet encoder as a foundation vision encoder and introduce DiffCut, an unsupervised zero-shot segmentation method that solely harnesses the output features from the final self-attention block. Through extensive experimentation, we demonstrate that using these diffusion features in a graph based segmentation algorithm, significantly outperforms previous state-of-the-art methods on zero-shot segmentation. Specifically, we leverage a recursive Normalized Cut algorithm that regulates the granularity of detected objects and produces well-defined segmentation maps that precisely capture intricate image details. Our work highlights the remarkably accurate semantic knowledge embedded within diffusion UNet encoders that could then serve as foundation vision encoders for downstream tasks. Project page: https://diffcut-segmentation.github.io
@@ 1 Introduction
In this work, we introduce DiffCut, a new method for zero-shot image segmentation which solely harnesses the encoder features of a pre-trained diffusion model in a recursive graph partitioning algorithm to produce fine-grained segmentation maps. Importantly, our method does not require any label from downstream segmentation datasets and its backbone has not been pre-trained on dense pixel annotations such as SAM [27]. We observe in Fig.˜1 that DiffCut produces sharp segments that nicely outline object boundaries. In comparison with the recent state-of-the-art unsupervised zero-shot segmentation method DiffSeg [1], the segments yielded by DiffCut, are better aligned with the semantic visual concepts, e.g. DiffCut is able to uncover the urban area as well as the boats in the middle row image. Our main contributions are as follows:
  • We leverage the features from the final self-attention block of a diffusion UNet encoder for the task of unsupervised image segmentation. In this context, we demonstrate that exploiting the inner patch-level alignment yields superior performance compared to merging self-attention maps as done in DiffSeg [1].
  • Compared to existing graph-based object localization methods e.g. TokenCut or MaskCut [30, 32], we push further and take advantage of a recursive Normalized Cut algorithm to generate dense segmentation maps. Via a partitioning threshold, the method is able to regulate the granularity of detected objects and consequently adapt the number of segments to the visual content.
@@ 3.1 Features Extraction
An input image is encoded into a latent via the VQ-encoder of the latent diffusion model and a small amount of gaussian noise is added to it (not shown in Fig.˜2). The obtained latent is passed to the diffusion UNet encoder, from which we only extract the output features, denoted $\hat{z}$, from its last self-attention block. This choice design has several motivations:
@@
In contrast to several methods that harness cross-attention modules for localization or segmentation tasks [39, 40, 41, 42], we deliberately choose not to depend on this mechanism. The accuracy of segmentation maps generated via attention modules heavily relies on the quality of the textual input, which often requires an automatic captioning model combined with a meticulous prompt design to reach competitive performance. Besides being constrained by the maximum number of input tokens, such approach is proved to be inaccurate in the presence of cohyponyms [64] and is prone to neglect subject tokens as the number of objects to detect becomes large [65]. The localization and segmentation capacity with a single forward inference is then constrained by the performance of the captioning model and the attention modules themselves. Exploiting only the intermediate diffusion features alleviate the computational cost of an additional captioning model and do not necessitate multiple forward inferences.
@@
Previous works [66, 67, 37] have shown that diffusion features provide precise semantic information shared across objects from different domains. Building on this observation, we hypothesize that the pyramidal architecture of the UNet encoder capture semantically rich image representations that are well-suited for zero-shot vision tasks. To validate this assumption, we exhibit the semantic coherence emerging in the UNet encoder, evidenced by a remarkable patch-level alignment in the output features of the final self-attention block. We in fact demonstrate that these features are sufficient to reach state-of-the-art zero-shot segmentation performance.
@@ 3.2 Recursive Feature Clustering
to find the eigenvector $x$ corresponding to the second smallest eigenvalue. In the ideal case, the clustering solution only takes two discrete values. Since the solution of Eq.˜2 is a continuous relaxation of the initial problem, $x$ contains continuous values and a splitting point has to be determined to partition it. To find the optimal partition, we examine $l$ evenly spaced points in $x$ and select the one resulting in the minimum NCut value.
@@
Based on our observations of the patch-level alignment of diffusion features illustrated in Fig.˜4, we assume that the normalized cut algorithm will produce sharp segments, each corresponding to a precise semantic concept as distinct objects would manifest as weakly connected components in a patch similarity matrix. Following this intuition, we construct an affinity matrix $\bm{W}$, by computing the cosine similarity, normalized between 0 and 1, between each pair of patches. As the NCut criterion evaluates both the dissimilarity between different segments and the similarity within each segment, we opt to emphasize inter-segments dissimilarity by raising each element to a positive integer power $\alpha$:
@@
Classical spectral clustering [68] requires setting a pre-defined number of clusters to partition the graph, which is a significant constraint in the context of zero-shot image segmentation where no prior knowledge on the visual content is available. We therefore adopt a recursive graph partitioning [55], which adapts the number of uncovered segments to the visual content via a threshold, denoted $\tau$, on the maximum partitioning cost. This hyperparameter stops the recursive partitioning of a segment when its NCut value exceeds it and thereby regulates the granularity of detected objects. We demonstrate that our soft thresholding process detailed above enhances the robustness of the method which delivers competitive performance across a wide range of $\tau$ values. This recursive clustering process is summarized in Supplementary˜A.
@@ 4 Experiments
Following existing works in image segmentation [21, 53, 52, 18], we use the following datasets for evaluation: a) Pascal VOC [70] (20 foreground classes), b) Pascal Context [71] (59 foreground classes), c) COCO-Object [72] (80 foreground classes), d) COCO-Stuff-27 merges the 80 things and 91 stuff categories in COCO-stuff into 27 mid-level categories, e) Cityscapes [73] (27 foreground classes) and f) ADE20K (150 foreground classes) [74]. An extra background class is considered in Pascal VOC, Pascal Context, and COCO-Object. We ignore their training sets and directly evaluate our method on the original validation sets, with the exception of COCO for which we evaluate on the validation split curated by prior works [21, 19].
@@
For all datasets, we report the mean intersection over union (mIoU), the most popular evaluation metric for semantic segmentation. Because our method does not provide a semantic label, we use the Hungarian matching algorithm [75] to assign the predicted masks to a ground truth mask. For datasets including a background class, we perform a many-to-one matching to the background label (Supplementary˜H). As in [1], we emphasize unsupervised adaptation (UA), language dependency (LD), and auxiliary image (AX). UA means that the specific method requires unsupervised training on the target dataset. Methods without the UA requirement are considered zero-shot. LD means that the method requires text input, such as a descriptive sentence for the image, to facilitate segmentation. AX means that the method requires an additional set of reference images.
@@
DiffCut builds on SSD-1B [35], a distilled version of Stable Diffusion XL [34]. The model takes an empty string as input and we set the timestep for denoising to $t=50$. To ensure a fair comparison when evaluating our method against baselines, we set a unique value for $\tau$ and $\alpha$ in all datasets, equal to $0.5$ and $10$ respectively. Following previous works, we make use of PAMR [69] to refine our segmentation masks. Our method runs on a single NVIDIA TITAN RTX (24GB) with input images of size $1024\times 1024$ and can segment an image in one second.
@@ 4.1 Results on Zero-shot Segmentation
Tab.˜1 reports the mIoU score for each baseline across the 6 benchmarks. Note that the numbers shown for COCO-Stuff and Cityscapes are taken from [1]. We complete ReCo [53] and MaskCLIP [52] scores with the results obtained in [50]. Other numbers are taken from [18]. We also note that DiffSeg tunes the sensible merging hyperparameter on a subset of images from the training set from the respective datasets. For a fair comparison, we evaluate the method fixing it to 1, as recommended in the original paper, and refine the masks obtained with PAMR. This baseline is denoted DiffSeg.
@@
Model LD AX UA VOC Context COCO-Object COCO-Stuff-27 Cityscapes ADE20K Extra-Training    IIC [19] 9.8 - - 6.7 6.4 -    MDC [76] - - - 9.8 7.1 -    PiCIE [21] - - - 13.8 12.3 -    PiCIE+H [21] - - - 14.4 - -    EAGLE [77] - - - 27.2 22.1 -    U2Seg [78] - - - 30.2 - -    STEGO [20] - - - 28.2 21.0 -    ACSeg [18] 53.9 - - 28.1 - - Training-free    ReCO [53] 25.1 19.9 15.7 26.3 19.3 11.2    MaskCLIP [52] 38.8 23.6 20.6 19.6 10.0 9.8    MaskCut $(k=5)$ [32] 53.8 43.4 30.1 41.7 18.7 35.7    DiffSeg [1] - - - 43.6 21.2 -    DiffSeg 49.8 48.8 23.2 44.2 16.8 37.7    DiffCut (Ours) 65.2 56.5 34.1 49.1 30.6 44.3
With our set of default hyperparameters, DiffCut significantly outperforms all other baselines despite not relying on language dependency, auxiliary images or unsupervised adaptation. On average, our method achieves a gain of +7.3 mIoU over the second best baseline. Notably, DiffCut exceeds MaskCut with an average improvement of +9.4 mIoU. Additionally, it outperforms the previous state-of-the-art method in unsupervised segmentation, DiffSeg, by +5.5 mIoU on COCO-Stuff and +9.4 mIoU on Cityscapes. The superiority of DiffCut over these two methods demonstrates our two key contributions: the high quality of our visual features for semantic segmentation and the flexibility of the recursive NCut algorithm in adjusting the number of segments according to the visual content of each image. The effectiveness of our method is further corroborated by our qualitative results shown in Fig.˜1. In comparison to DiffSeg, DiffCut provides finely detailed segmentation maps that more closely align with semantic concepts. Additional examples can be found in Supplementary˜N.
We note here that, as the granularity of annotations varies across target datasets, our fixed set of hyperparameters can not be in the optimal regime on each of them. Therefore, relaxing the condition on prior knowledge about the target dataset, we report in Supplementary˜G results of DiffCut where $\tau$ is loosely tuned using a small subset of annotated images from the target training split.
@@ 4.2 Semantic Coherence in Vision Encoders
As in [81], we collect patch representations from various vision encoders and store their corresponding target classes using the segmentation labels. Given $\hat{z}_{i}=\mathcal{E}(\mathbf{x_{1}})_{i}\in\mathbb{R}^{D_{v}}$ and $\hat{z}_{j}=\mathcal{E}(\mathbf{x_{2}})_{j}\in\mathbb{R}^{D_{v}}$, the patch representations of images $\mathbf{x_{1}}$ and $\mathbf{x_{2}}$ at respectively index $i$ and $j$, we compute their cosine similarity and use this score as a binary classifier to predict if the two patches belong to the same class. Given $l(\mathbf{x_{1}})_{i}$ and $l(\mathbf{x_{2}})_{j}$, the labels associated with the patches, if $l(\mathbf{x_{1}})_{i,j}=l(\mathbf{x_{2}})_{p,q}$, the target value for binary classification is 1, else 0. We present in Fig.˜3 the ROC curve and AUC score for our candidate models. We observe that SSD-1B UNet encoder [35] demonstrates greater patch-level alignment than any other candidate model with an AUC score of 0.83, even surpassing DINOv2 [5]. We also show the outstanding alignment between patch representations associated with semantically similar concepts with qualitative results in Fig.˜4. We provide additional qualitative examples of patch-level alignment in Supplementary˜M.
@@
A potential rationale for this observation lies in the superior semantic information retention of a diffusion model compared to alternative backbones, attributed to its inherent capacity to set a structural image layout, internally acquired during the training phase. These results provide insight into the strong clustering results presented in the previous section, as improved semantic coherence suggests that patches belonging to the same object are more effectively clustered.
@@ 4.3 Ablation study
DiffSeg proposes their own clustering algorithm based on a self-attention map merging process. As the original implementation uses a different diffusion backbone as ours, we validate the benefit of our method by swapping the original SDv1.4 with our stronger SSD-1B. For a fair comparison between methods, we use the default set of hyperparameters recommended in [1] and set the default merging threshold of DiffSeg to $0.5$ for all datasets. Tab.˜2 clearly validates the superiority of using rich semantic features in a recursive graph partitioning algorithm over the self-attention merging mechanism of DiffSeg. Qualitative results shown in Fig.˜1 further display the edge of DiffCut in uncovering semantic clusters. Shown results do not make use of the mask refinement module, explaining the gap with Tab.˜1.
@@
In DiffCut, the hyperparameter $\tau$ corresponds to the maximum graph partitioning cost allowed. In contrast, classical spectral clustering requires to explicitly set the number of segments to be found in the graph. To validate the benefit of the recursive approach over spectral clustering, we introduce a simple yet effective baseline dubbed AutoSC. [82] proposes a heuristic that estimates the number of connected components in a graph with the largest relative-eigen-gap between its Laplacian eigen-values. The larger the gap, the more confident the heuristic. In our context, the index of the eigen-value that maximizes this quantity can be interpreted as the number of clusters in an image. Thus, we define a set of exponents $\{1,5,10,15\}$ and determine the value $\alpha$ in this set such that its element-wise exponentiation of matrix $\bm{A}$ yields the largest Laplacian relative-eigen-gap. Then, we use the index of the eigen-value maximizing the gap as the number of clusters in a $k$-way spectral clustering performed with the algorithm proposed in [83]. As shown in Tab.˜2, DiffCut consistently outperforms AutoSC on all datasets, with a gain up to +3.5+3.5 on ADE20K, at the exception of COCO-Stuff where the latter yields slightly better results. Noting that AutoSC is already a state-of-the-art baseline on most benchmarks, this experiment confirms the relevance of the recursive Normalized Cut to uncover arbitrary numbers of segments.